# X-Vine Models for Multivariate Extremes

### v1

@@ B.1 Proofs for Section 2
First we state an auxiliary lemma linking the angular measure density $h$ to the exponent measure density $\lambda$ and the tail copula density $r$.
@@
The existence is clear: at every level $j\in\left\{{1,\ldots,d-1}\right\}$, the union $\bigcup_{e^{\prime}\in E_{j}}A_{e^{\prime}}$ of the complete unions is equal to $[d]$. Indeed, this holds for $E_{1}$ since $T_{1}$ is a tree on $[d]$ and it holds for the other edge sets by the recursive definition of the complete unions. The uniqueness is specific to the two nodes in the conditioned set of the single edge $e$ in $E_{d-1}$. Suppose that for some $a\in\mathcal{C}_{e}$ there is an edge set $E_{j}$, with $j\leqslant d-2$, for which there are two different edges, say $f$ and $g$, that both contain $a$ in their complete unions. If $f$ and $g$ are linked up by some edge $h$ in the next tree $T_{j+1}$, then $a\in D_{h}$ and thus, eventually $a\in D_{e}$ (see the paragraph following Definition 2.3), a contradiction. Otherwise, there exist two different edges $f^{\prime}$ and $g^{\prime}$ in $E_{j+1}$ that contain $a$ in their complete unions, and the argument can be repeated, if needed until the last tree has been reached. Finally, since the conditioning sets are nested and since $a$ does not belong to the conditioning set of $e$, it does not belong to the conditioning set of another edge either. ∎
@@
which is obviously true since $\gamma_{2}=1$ by assumption. Let $d\geqslant 4$ and suppose the formula is true for regular vine sequences of dimension up to $d-1$. Let $e=(a_{e},b_{e};D_{e})$ be the single edge in the deepest tree $T_{d-1}=(N_{d-1},E_{d-1})$ of the vine. By Lemma B.2, there exist unique edges $e_{j}\in E_{j}$ for $j\in\left\{{1,\ldots,d-1}\right\}$ with $e_{d-1}=e$ such that
@@
Moreover, the node $a_{e}$ belongs to the conditioned set of each edge $e_{j}$, so we can write $\mathcal{C}_{e_{j}}=\left\{{a_{e},b_{j}}\right\}$ for $j\in\left\{{1,\ldots,d-1}\right\}$ and some $b_{j}\in[d]\setminus\left\{{a_{e}}\right\}$. For $j\in\{2,\ldots,d-1\}$, we have $e_{j}=\left\{{e_{j-1},f_{j-1}}\right\}$ for some edge $f_{j-1}\in E_{j-1}$. The regular vine sequence $\mathcal{V}$ induces a regular vine sequence $\bar{\mathcal{V}}$ on $\bar{N}_{1}=[d]\setminus\left\{{a_{e}}\right\}$ with edge sets $\bar{E}_{j}=E_{j}\setminus\left\{{e_{j}}\right\}$ for $j\in\left\{{1,\ldots,d-2}\right\}$; this is the sub-vine induced by $\mathcal{V}$ on $[d]\setminus\left\{{a_{e}}\right\}$. By the induction hypothesis, we have
@@ B.3 Proofs for Section 4
This fact already implies that the simplifying assumption holds. To identify $c_{I;J}$, we can incorporate the constants $c(\alpha_{i},\rho)$ into the variables $x_{i}$ by marginal scale transformations without changing the copula. Hence, $c_{I;J}$ is equal to the copula density of the $|I|$-variate probability density function proportional to
@@ B.4 Proofs for Section 5
Proof of Theorem 5.1 and Remark 5.2.
@@
By Eq. (3.1) in Sklar’s theorem for tail copula densities (Proposition 3.1) applied to $I=\left\{{a_{e},b_{e}}\right\}$ and $J=D_{e}$, the last fraction is equal to $c_{a_{e},b_{e};D_{e}}\left({u_{a_{e}},u_{b_{e}};\bm{x}_{D_{e}}}\right)$ with $u_{a_{e}}=R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}})$ and $u_{b_{e}}=R_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}})$. Formula (5.2) in Remark 5.2 follows in the same way from the more general product formula (2.15) in Remark 2.5. ∎
@@
We have $e=(a_{e},b_{e};D_{e})=\left\{{f,g}\right\}$ with $f,g\in N_{d-1}=E_{d-2}$. The complete union of $e$ is $A_{e}=[d]$. Furthermore, upon switching the names of $f$ and $g$ if necessary, we have $A_{f}=D_{e}\cup\left\{{a_{e}}\right\}=[d]\setminus\left\{{b_{e}}\right\}$ and $A_{g}=D_{e}\cup\left\{{b_{e}}\right\}=[d]\setminus\left\{{a_{e}}\right\}$. We can apply the decomposition (5.1) to both $r$ and its $(d-1)$-dimensional marginal density $r_{A_{g}}$, as in the previous paragraph; see also Eq. (5.2). The two decompositions are the same, except that the one for $r$ has some additional factors, namely the ones that involve node $a_{e}$. For each tree, there is only one such additional factor, since in each tree $T_{j}$, there is only a single edge, say $e_{j}$, that contains $a_{e}$ in its complete union (Lemma B.2). For these edges $e_{j}$, we have in fact $a_{e}\in\mathcal{C}_{e_{j}}$, so we can write $e_{j}=(a_{e},b_{e_{j}};D_{e_{j}})$. Node that $e_{d-1}=e$ and $b_{e_{d-1}}=b_{e}$, while $e_{d-2}=f$. It follows that
@@
By Proposition C.1, this function is a valid tail copula density with margins $r_{12}$ and $r_{23}$. Furthermore,
@@
as there are $d-1$ edges in the first tree. Further, by the induction hypothesis and by scale equivariance (Lemma 3.2), we have $R_{a_{e}|D_{e}}(sx_{a_{e}}|s\bm{x}_{D_{e}})=R_{a_{e}|D_{e}}(sx_{a_{e}}|s\bm{x}_{D_{e}})$ and similarly for $a_{e}$ replaced by $b_{e}$. It follows that $r$ defined in (5.5) satisfies (2.5).
@@
$r(\bm{x})=r_{A_{g}}(\bm{x}_{A_{g}})\cdot r_{a_{e},b_{1}}(x_{a_{e}},x_{b_{1}})\cdot\prod_{j=2}^{d-1}c_{a_{e},b_{j};D_{e_{j}}}\left({R_{a_{e};D_{e_{j}}}(x_{a_{e}}|\bm{x}_{D_{e_{j}}}),R_{b_{j};D_{e_{j}}}(x_{b_{j}}|\bm{x}_{D_{e_{j}}})}\right).$

### v2

@@ B.1 Proofs for Section 2
Recall $\mathbb{E}=(0,\infty]^{d}\setminus\{\bm{\infty}\}$ for some dimension $d\geqslant 1$.
Lemma B.1.
Proof of Lemma B.1.
Given $R$, define the probability measure $P$ on $\mathbb{L}=\left\{{\bm{x}\in(0,\infty]^{d}:\min\bm{x}<1}\right\}$ by $P(B)=R(B\cap\mathbb{L})/R(\mathbb{L})$ for Borel sets $B$. (Note that $P$ is an inverted multivariate Pareto distribution as in Definition 6.1, although we do not use that fact in this proof.) For $j\in[d]$ and $x_{j}\in[0,1]$, we have, by (2.2),
$F_{j}(x_{j})=P(\{\bm{y}\in\mathbb{E}:y_{j}\leqslant x_{j}\})=x_{j}/R(\mathbb{L}).$
Note that $R(\mathbb{L})\geqslant 1$ by (2.2). For $u_{j}\in[0,1/R(\mathbb{L})]$, we have therefore
$F_{j}^{-1}(u_{j})=u_{j}\,R(\mathbb{L}).$
Let $C$ be the copula of $P$. For $\bm{u}\in[0,1/R(\mathbb{L})]^{d}$, we find, by Sklar’s theorem,
$C(\bm{u})=R(\bm{u}\,R(\mathbb{L}))/R(\mathbb{L})=R(\bm{u})$
in view of the homogeneity of $R$ in (2.3). The same equation then implies that $t^{-1}\cdot C(t\bm{x})=R(\bm{x})$ for all $\bm{x}\in(0,\infty)^{d}$ and all sufficiently small $t>0$. If some (but not all) coordinates of $\bm{x}$ are infinity, we repeat the same argument on the appropriate lower-dimensional margins of $C$. We conclude that (2.1) holds, so that $R$ is indeed a tail copula measure. ∎
By (2.3) and (2.9), the exponent measure $\Lambda$ is homogeneous of order $-1$ and thus determined by the angular measure $H$ on the unit simplex $\Delta_{d-1}=\left\{{\bm{x}\in[0,1]^{d}:\lVert{\bm{x}}\rVert_{1}=1}\right\}$ via
for $t>0$ and Borel sets $B\subseteq\Delta_{d-1}$, with $\lVert{\bm{x}}\rVert_{1}=\sum_{j=1}^{d}|x_{j}|$ for $\bm{x}\in\mathbb{R}^{d}$. Under the above smoothness assumption, the angular measure $H$ is concentrated on the relative interior of the unit simplex $\Delta_{d-1}$ and its density $h$ satisfies $h(\bm{w})=\lambda(\bm{w})$ for $\bm{w}\in\Delta_{d-1}$ (Dombry et al., 2016). In Lemma B.2 below, we give a proof of this fact and link the angular measure density $h$ to the tail copula density $r$.
@@
Properties of regular vine sequences.
Recall Definition 2.3 on the complete union $A_{e}$, conditioning set $D_{e}$ and conditioned sets $\mathcal{C}_{e}=\mathcal{C}_{e,a}\cup\mathcal{C}_{e,b}$ of an edge $e\in\bigcup_{j=2}^{d}E_{j}$ in a regular vine $\mathcal{V}=(T_{1},\ldots,T_{d-1})$ on $d$ elements. If $j\geqslant 2$, then $D_{a}\cup D_{b}\subseteq D_{e}$; indeed, by proximity, $a\cap b=\left\{{k}\right\}$ with $k\in N_{j-1}$, and $D_{a}\subset A_{k}\subset A_{a}\cap A_{b}=D_{e}$; similarly $D_{b}\subset D_{e}$. Further, the conditioned sets are related by $\mathcal{C}_{e,a}=\mathcal{C}_{e}\cap\mathcal{C}_{a}$ and $\mathcal{C}_{e,b}=\mathcal{C}_{e}\cap\mathcal{C}_{b}$. To see this, note that, by proximity, we have $a=\left\{{a_{1},k}\right\}$ and $b=\left\{{b_{1},k}\right\}$ for three distinct elements $a_{1},b_{1},k\in N_{j-1}$, and $\mathcal{C}_{e,a}=A_{a}\setminus A_{b}=\left({A_{a_{1}}\cup A_{k}}\right)\setminus\left({A_{b_{1}}\cup A_{k}}\right)\subseteq A_{a_{1}}\setminus A_{k}=\mathcal{C}_{a,a_{1}}$. Since both the left- and right-hand side are singletons, we must have $\mathcal{C}_{e,a}=\mathcal{C}_{a,a_{1}}=\mathcal{C}_{e}\cap\mathcal{C}_{a}$; similarly $\mathcal{C}_{e,b}=\mathcal{C}_{b,b_{1}}=\mathcal{C}_{e}\cap\mathcal{C}_{b}$.
@@
The existence is clear: at every level $j\in\left\{{1,\ldots,d-1}\right\}$, the union $\bigcup_{e^{\prime}\in E_{j}}A_{e^{\prime}}$ of the complete unions is equal to $[d]$. Indeed, this holds for $E_{1}$ since $T_{1}$ is a tree on $[d]$ and it holds for the other edge sets by the recursive definition of the complete unions.
The uniqueness is specific to the two nodes in the conditioned set of the single edge $e$ in $E_{d-1}$. Suppose that for some $a\in\mathcal{C}_{e}$ there is an edge set $E_{j}$, with $j\leqslant d-2$, for which there are two different edges, say $f$ and $g$, that both contain $a$ in their complete unions. If $f$ and $g$ are linked up by some edge $h$ in the next tree $T_{j+1}$, then $a\in D_{h}$ and thus, eventually $a\in D_{e}$ (see the paragraph following Definition 2.3), a contradiction. Otherwise, there exist two different edges $f^{\prime}$ and $g^{\prime}$ in $E_{j+1}$ that contain $a$ in their complete unions, and the argument can be repeated, if needed until the last tree has been reached. Finally, since the conditioning sets are nested and since $a$ does not belong to the conditioning set of $e$, it does not belong to the conditioning set of another edge either. ∎
@@
which is obviously true since $\gamma_{2}=1$ by assumption.
Let $d\geqslant 4$ and suppose the formula is true for regular vine sequences of dimension up to $d-1$. Let $e=(a_{e},b_{e};D_{e})$ be the single edge in the deepest tree $T_{d-1}=(N_{d-1},E_{d-1})$ of the vine. By Lemma B.3, there exist unique edges $e_{j}\in E_{j}$ for $j\in\left\{{1,\ldots,d-1}\right\}$ with $e_{d-1}=e$ such that
@@
Moreover, the node $a_{e}$ belongs to the conditioned set of each edge $e_{j}$, so we can write $\mathcal{C}_{e_{j}}=\left\{{a_{e},b_{j}}\right\}$ for $j\in\left\{{1,\ldots,d-1}\right\}$ and some $b_{j}\in[d]\setminus\left\{{a_{e}}\right\}$. In particular, $a_{e}$ is a leaf node in $T_{1}$. For $j\in\{2,\ldots,d-1\}$, we have $e_{j}=\left\{{e_{j-1},f_{j-1}}\right\}$ for some edge $f_{j-1}\in E_{j-1}$.
Thanks to the stated properties of $a_{e}$, the regular vine sequence $\mathcal{V}$ induces a regular vine sequence $\bar{\mathcal{V}}$ on $\bar{N}_{1}=[d]\setminus\left\{{a_{e}}\right\}$ with edge sets $\bar{E}_{j}=E_{j}\setminus\left\{{e_{j}}\right\}$ for $j\in\left\{{1,\ldots,d-2}\right\}$; this is the sub-vine induced by $\mathcal{V}$ on $[d]\setminus\left\{{a_{e}}\right\}$. By the induction hypothesis, we have
@@
Remark B.4 (Vine telescoping product on sub-vines).
$\gamma_{A_{f}}=\prod_{e\in E_{1}:A_{e}\subseteq A_{f}}\gamma_{e}\cdot\prod_{i=2}^{j}\prod_{e\in E_{i}:A_{e}\subseteq A_{f}}\frac{\gamma_{\{a_{e},b_{e}\}|D_{e}}}{\gamma_{a_{e}|D_{e}}\cdot\gamma_{b_{e}|D_{e}}}.$ (B.1)
Remark B.5 (Vine telescoping product when $\gamma_{j}$ is not necessarily $1$.).
Another generalisation of the vine telescoping product in Lemma 2.4 concerns removing the assumption that $\gamma_{j}=1$ for all $j\in[d]$. If $\gamma_{J}\in(0,\infty)$ for every non-empty $J\subseteq[d]$ but the scalars $\gamma_{j}$ are not necessarily equal to one, the vine telescoping product in Lemma 2.4 becomes
$\gamma_{[d]}=\prod_{j\in[d]}\gamma_{j}^{1-\deg(j)}\cdot\prod_{e\in E_{1}}\gamma_{e}\cdot\prod_{j=2}^{d-1}\prod_{e=\left\{{a,b}\right\}\in E_{j}}\frac{\gamma_{D_{e}}\cdot\gamma_{A_{e}}}{\gamma_{A_{a}}\cdot\gamma_{A_{b}}},$ (B.2)
where $\deg(j)$ is the degree of $j$ in $T_{1}$, that is, the number of edges in that tree to which it belongs, or, equivalently, its number of neighbours in that tree. To obtain (B.2), apply (2.13) to the coefficients $\gamma_{J}^{\prime}=\gamma_{J}/\prod_{j\in J}\gamma_{j}$ for non-empty $J\subseteq[d]$ and simplify.
@@ B.2 Proofs for Section 3
Proof of Proposition 3.5.
First, we show that $r$ is a tail copula density. To this end, we need to show two properties: the normalization constraint (2.4) and the homogeneity relation (2.5).
To show the homogeneity relation (2.5), let $t\in(0,\infty)$, and let $\bm{x}\in(0,\infty)^{d}$ be such that $r_{J}(\bm{x}_{J})>0$. Then
$(1-|J|)(1-|I|)-|I|\cdot|J|=1-|I|-|J|=1-d.$
To show the normalization constraint (2.4), let $k\in\left\{{1,\ldots,d}\right\}$ and $x_{k}\in(0,\infty)$. We need to show that $\int_{(0,\infty)^{d-1}}r(\bm{x})\,\mathrm{d}\bm{x}_{\setminus k}=1$, where the integral is over all coordinates except for the $k$th one. There are two cases: $k\in I$ or $k\in J$.
First, if $k\in I$, the integral of interest is
as $r_{k\cup J}$ is a tail copula density.
Second, if $k\in J$, the integral of interest is
since $r_{J}$ is a tail copula density.
The proof that $r$ has the required margins $r_{i\cup J}$ for $i\in I$ is similar to the case $k\in I$ in the previous part of the proof.
Finally, statement (3.1) easily follows by applying Sklar’s theorem to $r_{I|J}(\,\cdot\,|\bm{x}_{J})$. ∎
Remark B.6.
In the special case that $J$ is a singleton, $J=\left\{{j}\right\}$, we have $I=[d]\setminus j$, the $(d-1)$ tail copula densities $(r_{i,j})_{i\in[d]\setminus j}$ are bivariate, and the common-margin assumption is automatic by (2.2); in particular, $r_{i|j}(x_{i}|x_{j})=r_{i,j}(x_{i},x_{j})$. Further, the scale-invariance property (3.2) enforces the simplifying assumption, so that the construction (3.3) becomes
@@ B.3 Proofs for Section 4
This fact already implies that the simplifying assumption holds.
To identify $c_{I;J}$, we can incorporate the constants $c(\alpha_{i},\rho)$ into the variables $x_{i}$ by marginal scale transformations without changing the copula. Hence, $c_{I;J}$ is equal to the copula density of the $|I|$-variate probability density function proportional to
@@ B.4 Proofs for Section 5
Proof of Theorem 5.1 and Remark B.7.
@@
By Eq. (3.1) in Sklar’s theorem for tail copula densities (Proposition 3.1) applied to $I=\left\{{a_{e},b_{e}}\right\}$ and $J=D_{e}$, the last fraction is equal to $c_{a_{e},b_{e};D_{e}}\left({u_{a_{e}},u_{b_{e}};\bm{x}_{D_{e}}}\right)$ with $u_{a_{e}}=R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}})$ and $u_{b_{e}}=R_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}})$. Formula (B.4) in Remark B.7 follows in the same way from the more general product formula (B.1) in Remark B.4. ∎
Remark B.7.
The decomposition (5.1) is based on the vine telescoping product formulas (2.13)–(2.14) in Lemma 2.4. By applying the more general formula (B.1) in Remark B.4, we find, for every edge $f\in E_{j}$ with $j\in\left\{{1,\ldots,d-1}\right\}$, a decomposition of the marginal tail copula density on the complete union $J=A_{f}\subseteq[d]$, with cardinality $|J|=j+1$: for $\bm{x}_{J}\in(0,\infty)^{J}$, we have
$r_{J}(\bm{x}_{J})=\prod_{\begin{subarray}{c}e\in E_{1}\\ A_{e}\subseteq J\end{subarray}}r_{a_{e},b_{e}}(x_{a_{e}},x_{b_{e}})\cdot\prod_{i=2}^{|J|-1}\prod_{\begin{subarray}{c}e\in E_{i}\\ A_{e}\subseteq J\end{subarray}}c_{a_{e},b_{e};D_{e}}\left({R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}}),R_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}});\bm{x}_{D_{e}}}\right).$ (B.4)
The decomposition (B.4) is true only for subsets $J\subseteq[d]$ of the form $J=A_{f}$ for some edge $f$ in the regular vine sequence.
@@
We have $e=(a_{e},b_{e};D_{e})=\left\{{f,g}\right\}$ with $f,g\in N_{d-1}=E_{d-2}$. The complete union of $e$ is $A_{e}=[d]$. Furthermore, upon switching the names of $f$ and $g$ if necessary, we have $A_{f}=D_{e}\cup\left\{{a_{e}}\right\}=[d]\setminus\left\{{b_{e}}\right\}$ and $A_{g}=D_{e}\cup\left\{{b_{e}}\right\}=[d]\setminus\left\{{a_{e}}\right\}$. We can apply the decomposition (5.1) to both $r$ and its $(d-1)$-dimensional marginal density $r_{A_{g}}$, as in the previous paragraph; see also Eq. (B.4). The two decompositions are the same, except that the one for $r$ has some additional factors, namely the ones that involve node $a_{e}$. For each tree, there is only one such additional factor, since in each tree $T_{j}$, there is only a single edge, say $e_{j}$, that contains $a_{e}$ in its complete union (Lemma B.3). For these edges $e_{j}$, we have in fact $a_{e}\in\mathcal{C}_{e_{j}}$, so we can write $e_{j}=(a_{e},b_{e_{j}};D_{e_{j}})$. Node that $e_{d-1}=e$ and $b_{e_{d-1}}=b_{e}$, while $e_{d-2}=f$. It follows that
@@
By Proposition 3.5, this function is a valid tail copula density with margins $r_{12}$ and $r_{23}$. Furthermore,
@@
as there are $d-1$ edges in the first tree. Further, by the induction hypothesis and by scale equivariance (Lemma 3.2), we have $R_{a_{e}|D_{e}}(sx_{a_{e}}|s\bm{x}_{D_{e}})=R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}})$ and similarly for $a_{e}$ replaced by $b_{e}$. It follows that $r$ defined in (5.6) satisfies (2.5).
@@
$r(\bm{x})=r_{A_{g}}(\bm{x}_{A_{g}})\cdot r_{a_{e},b_{1}}(x_{a_{e}},x_{b_{1}})\cdot\prod_{j=2}^{d-1}c_{a_{e},b_{j};D_{e_{j}}}\left({R_{a_{e}|D_{e_{j}}}(x_{a_{e}}|\bm{x}_{D_{e_{j}}}),R_{b_{j}|D_{e_{j}}}(x_{b_{j}}|\bm{x}_{D_{e_{j}}})}\right).$
@@
Proof of Proposition 5.8.
Let $\bm{x}\in(0,\infty)^{d}$. As there $d-1$ edges in the first tree, we have
The first product on the right-hand side converges to $\prod_{e\in E_{1}}r_{a_{e},b_{e}}(x_{a_{e}},x_{b_{e}})$ as $t\to 0$ by Assumption 5.7(i). It remains to show that for every $j\in\left\{{2,\ldots,d-1}\right\}$, every $e\in E_{j}$ and every $i\in\mathcal{C}_{e}=\left\{{a_{e},b_{e}}\right\}$, we have
$\lim_{t\searrow 0}C_{i|D_{e}}\left({tx_{i}|t\bm{x}_{D_{e}}}\right)=R_{i|D_{e}}\left({x_{i}|\bm{x}_{D_{e}}}\right).$ (B.6)
By properties of regular vine sequences as explained in Section 2.2, there exists an edge $f=(a_{f},b_{f};D_{f})\in E_{j-1}$ such that $f\in e$ and $i\in\mathcal{C}_{f}=\left\{{a_{f},b_{f}}\right\}$ and $A_{f}=D_{e}\cup\left\{{i}\right\}$. The proof of (B.6) is by recursion on $j$.
If $j=2$, then $f\in E_{1}$ so that $D_{f}=\varnothing$ while $D_{e}$ is a singleton, say $D_{e}=\left\{{d_{e}}\right\}$, and thus $f=\left\{{i,d_{e}}\right\}$. But then (B.6) follows from Assumption 5.7(i) and Lemma B.8 below.
Let $j\geqslant 3$ and assume (B.6) has been shown for edges in $E_{1}\cup\cdots\cup E_{j-1}$. Let $k$ be the other node in $\mathcal{C}_{f}$ but $i$, that is, $\left\{{a_{f},b_{f}}\right\}=\mathcal{C}_{f}=\left\{{i,k}\right\}$. Then by (5.9) applied to $f$, we have
and thus
By the induction hypothesis, the two arguments of $C_{i|k;D_{f}}$ converge to $R_{i|D_{f}}(x_{i}|\bm{x}_{D_{f}})$ and $R_{k|D_{f}}(x_{k}|\bm{x}_{D_{f}})$, respectively. As $f\in E_{j-1}$ and $j\geqslant 3$, the pair copula density $c_{i,k;D_{f}}$ is continuous by Assumption 5.7(ii). In view of Corollary B.10 below, $C_{i|k;D_{f}}$ is continuous in its two arguments too, and thus
By the recursive property (5.7), the last expression is equal to $R_{i|D_{f}\cup k}\left({x_{i}|\bm{x}_{D_{f}\cup k}}\right)$. As $D_{f}\cup k=D_{e}$, the convergence (B.6) follows. ∎
Lemma B.8.
The copula $C$ of $c$ and the tail copula $R$ of $r$ then satisfy
Proof of Lemma B.8.
But then
Lemma B.9.
Proof of Lemma B.9.
$\lim_{y^{\prime}\to y}\int_{I}\left|{f_{y^{\prime}}(x)-f_{y}(x)}\right|\,\mathrm{d}x=0.$ (B.7)
The first term on the right-hand side converges to zero by (B.7) and the second term vanishes asymptotically by the dominated convergence theorem. ∎
Corollary B.10.
Proof of Corollary B.10.
Apply Lemma B.9 to $f=c$ and $I=J=(0,1)$. ∎